OpenStudy (anonymous):

Line AB passes through A(-3, 0) and B(-6, 5). What is the equation of the line that passes through the origin and is parallel to line AB?

OpenStudy (anonymous):

5x-3y=0 -x+3y=0 -5x-3y=0 3x+5y=0 -3x+5y=0

OpenStudy (anonymous):

start by using slope formula to find slope the use the point slope fromula to find the funtion

OpenStudy (anonymous):

It's one of those I've tried I cannot figure it out.

OpenStudy (anonymous):

what is the slope between those points

OpenStudy (anonymous):

3,5

OpenStudy (anonymous):

you find slope by this formula y1-y2/x1-x2 find the slope then we'll go from there

OpenStudy (anonymous):

I think it's 3x+5y=0

OpenStudy (anonymous):

i don't the slope is -3\5 the we will but that in to the point slope formula

OpenStudy (anonymous):

so you shold have the funtion y=-3/5x+1.4

OpenStudy (anonymous):

dose that work